5x2 +16x +3 =0
I only need to solve this in QF and don't need to find the discriminant.

Answer by jst_kidding(6)

5x^2 + 16x + 3=0
5x^2 + 15x + x + 3=0
5x(x+3)+1(x+3)=0
(5x+1)(x+3)=0
therefore,
x= -1/5 and -3